OpenStudy (anonymous):

the quantity 7+3 radical 2 times the quantity 9 minus 5 radical 2.

OpenStudy (campbell_st):

\[(7+3\sqrt{2})\times(9 - 5\sqrt{2})\] this is a binomial product problem \[7\times9 - 7\times5\sqrt{2} + 9\times 3\sqrt{2} + 3\times-5\times \sqrt{2}\times \sqrt{2}\] simplifying gives \[63 - 8\sqrt{2} - 30\] final answer \[33 - 8\sqrt{2}\]

OpenStudy (anonymous):

thank you very much :)